OpenStudy (anonymous):

a car moves at 60 km/hr along straight highway how many 80 seconds will it take for the car to travel a distance of 10 km?

OpenStudy (anonymous):

How many intervals of 80 seconds? Well, we know that \[ \frac{10}{60}=\frac{1}{6} \text{ hours} \]And we also know: \[3600\text{ sec}=1\text{ hr}\]So, the ride will take\[ \frac{3600}{60}=60\text{ sec}=\frac{1}{80}\cdot60\text{ (80-second intervals)}=.75\text{ (80-second intervals)} \] And we are done.

OpenStudy (anonymous):

In an hour, the car travels 60 km. So, in one sixth of an hour, the car travels 10 km. That's 10 km in 10 minutes. Multiply minutes by 60 to get seconds and you're left with 600s, or 7.5 80s intervals

OpenStudy (anonymous):

Oh, sh*ite, yet again, I screwed up... Ahh, it's 7.5 like RyanE stated...